OpenStudy (anonymous):

how do you solve for cubed factoring example 8y^3-27x^3

OpenStudy (anonymous):

Just cube root the whole function. \[\sqrt[3]{8y-27x}\] Then simpify 8 and 27 by taking the cube root of that too. Got it?

OpenStudy (anonymous):

\(a^3-b^2=(a-b)(a^2+ab+b^2)\) with \(a=2y,b=3x\) in other words you are supposed to recognize both terms as perfect cubes

OpenStudy (anonymous):

sorry to say you cannot "cube root" the whole thing. your job is to factor. hope it is clear

OpenStudy (anonymous):

oh except i made a typo, it should be \[a^3-b^3=(a-b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

Well you pretty much get the same exact answer because they are perfect cubes.

OpenStudy (anonymous):

thanks so much this will help for tomorrows quiz !!

OpenStudy (anonymous):

not to be picky, but they are not at all the same thing asking to factor is not the same as asking to take the cube root.

OpenStudy (anonymous):

But wouldn't the formula you wrote one that could be used to factor out a cubed root ?